OpenStudy (anonymous):

Solution of this equation sqrt x+10-9=-7

OpenStudy (jjuden):

- 6 is the solution.

OpenStudy (anonymous):

how did you get that?

OpenStudy (jjuden):

√ (x + 10) - 9 = - 7 √ (x + 10) = 2 x + 10 = 4 x = - 6

OpenStudy (anonymous):

okay can you help with another?

OpenStudy (jjuden):

ok

OpenStudy (anonymous):

4(3-x)^4/3-5=59

OpenStudy (jjuden):

4(3-x)^4/3 -5 = 59 4(3 - x)4/3 = 64 (3 - x)4/3 = 16 (3 - x)2/3 = 4 ³√(3 - x) = 2 (3 - x) = 8 x = -5

OpenStudy (anonymous):

wow thank you so much!

OpenStudy (jjuden):

your welcome :)